OpenStudy (anonymous):

Which is equivalent to i224? 1, -1, 0 or i?

OpenStudy (anonymous):

i^(224)= (i^(4))^(56)= 1^56=1

OpenStudy (anonymous):

1 is the answer

OpenStudy (anonymous):